# Optimal Power Flow Pursuit via Feedback-based Safe Gradient Flow

### v2

@@ III-A Feedback-based online algorithm
In the next section, we analyze the convergence and stability properties of the proposed feedback-based SGF (9).
Figure 1: Closed-loop implementation of the proposed online feedback optimization algorithm. The proposed scheme leverages measurements of voltages and pseudo-measurements (if available) to solve the AC OPF problem (4).
III-B Stability analysis and constraint satisfaction guarantees
@@
Lemma III.4 (Lipschitz continuity).
@@
Lemma III.4 follows from [35, Theorem 3.6], and by the compactness of the sets $\mathcal{E}_{M}$ and $\mathcal{E}_{J}$. This result ensures existence and uniqueness of solutions for the proposed feedback-based safe gradient flow [36, Ch. 3].
@@
Theorem III.5 (Practical local exponential stability).
@@
Lemma III.6 (Practical forward invariance).
Let the conditions in Theorem III.5 be satisfied, and let $\boldsymbol{u}(t)$, $t\geq t_{0}$, be the unique trajectory of (9), and $\boldsymbol{\nu}(t)$ be the corresponding voltage magnitudes. Define the set
@@
The proof is provided in Appendix -C. Lemma III.6 establishes forward invariance of a set $\mathcal{F}_{s}$, which is a subset of $\mathcal{F}_{e}$ and an inflation of $\mathcal{F}$ (more details about $\mathcal{F}_{s}$ are provided in the proof); clearly, $\mathcal{F}_{e}$ tends to the set $\mathcal{F}$ with the decreasing of the error in the computation of the Jacobian and the measurement errors, which implies that $\mathcal{F}_{s}$ tends to $\mathcal{F}$ too. If these errors are small, the voltage violation is practically negligible.
@@
Corollary III.7 (Error-free implementation).
Let all the conditions in Theorem III.5 be satisfied, and assume that there are no measurement errors, i.e., $E_{M}=0$, and no errors in the Jacobian, i.e., $E_{J}=0$ and. Let $\boldsymbol{u}(t)$, $t\geq t_{0}$, be the unique trajectory of (7). Then it holds that
@@
Lemma III.8 (Forward invariance in error-free implementation).
Let the conditions in Theorem III.5 be satisfied, and assume that there are no measurement errors (i.e., $E_{M}=0$). Let $\boldsymbol{u}(t)$, $t\geq t_{0}$, be the unique trajectory of (7), and $\boldsymbol{\nu}(t)$ be the corresponding trajectory of the voltages. Then, (7) renders the set $\mathcal{F}$ forward-invariant. In particular: (i) if $\nu_{i}(t_{0})\in[\underline{V},\bar{V}]$, then $\nu_{i}(t)\in[\underline{V},\bar{V}]$ for all $t\geq t_{0}$; (ii) if $\nu_{i}(t_{0})\notin[\underline{V},\bar{V}]$, then there exists $t^{\prime}\geq t_{0}$ such that $\nu_{i}(t)\in[\underline{V},\bar{V}]$ for all $t\geq t^{\prime}$. $\Box$
Corollary III.7 quantifies the error in the convergence to $\boldsymbol{u}^{*}$, and certify local exponential stability properties for the proposed method. Lemma III.8 establishes that, for the case with no measurement errors and with the exact computation of the Jacobian matrix, the proposed method ensures that voltages are satisfied anytime.
III-C Extension to virtual power plants
In this section, we introduce an extension of our formulation for virtual power plants (VPPs). In this case, the goal is to coordinate the operation of the DERs to regulate voltages and to provide ancillary services to the bulk power system. In particular, the coordination is to ensure that the active and reactive powers at the substation track a reference $\{P_{0,set},Q_{0,set}\}$.
@@ IV-A Simulation setup
Finally, we consider the following cost function for the SGF, PDM, and BO:
$C_{P}(\boldsymbol{u})=\sum_{i\in{\mathcal{G}}}c_{p}\left(\frac{s_{n,i}-p_{i}}{s_{n,i}}\right)^{2}+c_{q}\left(\frac{q_{i}}{s_{n,i}}\right)^{2}$ (22)
in an effort to minimize active power curtailment and inverter power losses, with $c_{p}=3$ and $c_{q}=1$.
Volt/Var control and overvoltage protection scheme
@@
Figure 4: (a) Overvoltages for NC. (b) Overvoltages for VVC. (c) Overvoltages for PDM. (d) Overvoltages for SGF. (e) Overvoltage duration times. (f) Linear map error: $||\hat{H}(\boldsymbol{u};\boldsymbol{p}_{l},\boldsymbol{q}_{l})-H(\boldsymbol{u};\boldsymbol{p}_{l},\boldsymbol{q}_{l})||^{2}$, where $\boldsymbol{u}$ is picked from the SGF algorithm.
@@ IV-B Results
In Fig. 4, one can see the maximum voltage observed at every time step, as well as the number of impacted nodes where we observed a voltage greater than $\bar{V}$. It can be seen that only the SGF method does not lead to voltage violations. VVC performs well, although as shown hereafter, it leads to larger system losses and a greater cumulative cost than SGF or PDM. The overall voltage profile is also shifted downward due to its proportional feedback control. One can observe the spikes in the NC method due to multiple disconnections of DERs because of a prolonged overvoltage duration. Finally, one can see that with PDM, the voltages oscillate around the threshold voltage of 1.05 pu.
@@
Finally, Fig. 4(f) shows the error between the linear approximation of the power flow equations and the non-linear power flow equations, validating our choice for the linear map.
@@ V Conclusions
By [23, Lemma 5.11 and Theorem 5.6(iii)], $\bar{F}_{\beta}(\boldsymbol{u},\mymathbb{0},\mymathbb{0})$ is differentiable at $\boldsymbol{u}^{*}$ and its Jacobian $E=\frac{\partial\bar{F}_{\beta}(\boldsymbol{u},\mymathbb{0},\mymathbb{0})}{\partial\boldsymbol{u}}\mid_{\boldsymbol{u}=\boldsymbol{u}^{*}}$ is negative definite. Recall that $e_{1}=-\lambda_{\max}(E)$ and $e_{2}=-\lambda_{\min}(E)$. Let $P:=\int_{0}^{\infty}(\exp(E\zeta)^{\top}\exp(E\zeta)d\zeta$, and then by [36, Theorem 4.12], it holds that $PE+E^{\top}P=-\mathbf{I}_{n}$, and $\frac{1}{2e_{2}}\|u-u^{*}\|_{2}^{2}\leq(u-u^{*})^{\top}P(u-u^{*})\leq\frac{1}{2e_{1}}\|u-u^{*}\|_{2}^{2}$. Let $V_{1}(\boldsymbol{u}):=$$(\boldsymbol{u}-\boldsymbol{u}^{*})^{\top}P(\boldsymbol{u}-\boldsymbol{u}^{*})$; then we bound $2(\boldsymbol{u}-\boldsymbol{u}^{*})^{\top}P\bar{F}_{\beta}(\boldsymbol{u},\mymathbb{0},\mymathbb{0})$ and then leverage this bound to estimate $\dot{V}_{1}$:
@@
In addition, we note that that for any $a\geq 0,~{}b>0$, $y(t)=y(t_{0})\exp{(-b(t-t_{0}))}+\frac{a}{b}(1-\exp{(-b(t-t_{0})})$ is the solution of $\dot{y}=-by+a,y(t_{0})=y(t_{0})$. Hence by the Comparison Lemma [36, Lemma 3.4], it follows that

### v3

@@ III-A Feedback-based online algorithm
III-B Remarks on the linear model
In the following, we comment on the linear approximation used in this paper. We start with the power flow equations (2) that we linearize around a given voltage profile $\bar{\boldsymbol{v}}=[\bar{v}_{1},...,\bar{v}_{N}]^{\top}$. Let us consider $\boldsymbol{d}\in\mathbb{R}^{N}$ capturing the deviations around the linearization point. We have:
$\boldsymbol{s}=\text{diag}(\bar{\boldsymbol{v}}+\boldsymbol{d})\left(\boldsymbol{\bar{y}}^{*}v_{0}^{*}+\boldsymbol{Y}^{*}\bar{\boldsymbol{v}}^{*}\right)+\text{diag}(\boldsymbol{d})\left(\boldsymbol{Y}^{*}\bar{\boldsymbol{v}}^{*}\right)+\\ \text{diag}(\boldsymbol{d})\left(\boldsymbol{Y}^{*}\boldsymbol{d}^{*}\right).$ (13)
If we discard the second-order terms $\text{diag}(\boldsymbol{d})\left(\boldsymbol{Y}^{*}\boldsymbol{d}^{*}\right)$, and considering the following choice for the nominal voltage profile:
$\bar{\boldsymbol{v}}=-\boldsymbol{Y}^{-1}\boldsymbol{\bar{y}}v_{0},$ (14)
equation (13) becomes:
$\text{diag}(\bar{\boldsymbol{v}}^{*})\boldsymbol{Y}\boldsymbol{d}=\boldsymbol{s}^{*}.$ (15)
Let $\bar{\boldsymbol{\rho}}\in\mathbb{R}^{N}$ be the vector collecting the magnitudes of voltages $\bar{\boldsymbol{v}}$, and define $\bar{\boldsymbol{a}}:=\{\cos(\bar{\theta}_{n})\}_{n\in\mathcal{N}}$, $\bar{\boldsymbol{b}}:=\{\sin(\bar{\theta}_{n})\}_{n\in\mathcal{N}}$ with $\bar{\theta}_{i}$ the angle of the nominal voltage $\bar{v}_{i}$. A solution of (15) can be expressed as $\boldsymbol{d}=\boldsymbol{Y}^{-1}\text{diag}^{-1}(\bar{\boldsymbol{v}}^{*})\boldsymbol{s}^{*}$. Expanding this expression, and defining matrices:
$\bar{\boldsymbol{R}}=\boldsymbol{Z}_{R}\text{diag}(\bar{\boldsymbol{a}})(\text{diag}(\bar{\boldsymbol{\rho}}))^{-1}-\boldsymbol{Z}_{I}\text{diag}(\bar{\boldsymbol{b}})(\text{diag}(\bar{\boldsymbol{\rho}}))^{-1}$ (16)
$\bar{\boldsymbol{B}}=\boldsymbol{Z}_{I}\text{diag}(\bar{\boldsymbol{a}})(\text{diag}(\bar{\boldsymbol{\rho}}))^{-1}+\boldsymbol{Z}_{R}\text{diag}(\bar{\boldsymbol{b}})(\text{diag}(\bar{\boldsymbol{\rho}}))^{-1},$
where $\boldsymbol{Z}_{R}:=\Re\{\boldsymbol{Y}^{-1}\}$ and $\boldsymbol{Z}_{I}:=\Im\{\boldsymbol{Y}^{-1}\}$, one can write:
$\boldsymbol{v}\approx\left(\bar{\boldsymbol{R}}+j\bar{\boldsymbol{B}}\right)\boldsymbol{p}_{\mathrm{net}}+\left(\bar{\boldsymbol{B}}-j\bar{\boldsymbol{R}}\right)\boldsymbol{q}_{\mathrm{net}}+\bar{\boldsymbol{v}}.$ (17)
If the entries of $\bar{\boldsymbol{v}}$ dominate those in $\boldsymbol{d}$, then $\boldsymbol{\rho}+\Re\{\boldsymbol{d}\}$ serves as a first order approximation for the voltage magnitudes. Thus, one can write:
$\hat{H}(\boldsymbol{u};\boldsymbol{p}_{l},\boldsymbol{q}_{l}):=\bar{\boldsymbol{R}}(\boldsymbol{\Gamma}_{R}\boldsymbol{u}+\boldsymbol{p}_{l})+\bar{\boldsymbol{B}}(\boldsymbol{\Gamma}_{B}\boldsymbol{u}+\boldsymbol{q}_{l})+\bar{\boldsymbol{\rho}},$ (18)
with $\boldsymbol{\Gamma}_{R}\in\mathbb{R}^{N\times 2G}$ and $\boldsymbol{\Gamma}_{B}\in\mathbb{R}^{N\times 2G}$ matrices filled with 0 and 1 such that $\boldsymbol{\Gamma}_{R}\boldsymbol{u}+\boldsymbol{p}_{l}=\boldsymbol{p}_{\mathrm{net}}$ and $\boldsymbol{\Gamma}_{B}\boldsymbol{u}+\boldsymbol{q}_{l}=\boldsymbol{q}_{\mathrm{net}}$. Notice that equation (18) can be written as in (11). Effectively, the approximate Jacobian $J_{\hat{H}}=\bar{\boldsymbol{R}}\boldsymbol{\Gamma}_{R}+\bar{\boldsymbol{B}}\boldsymbol{\Gamma}_{B}$ no longer depends on $\boldsymbol{u}$ and the non-controllable powers $\boldsymbol{p}_{l},\boldsymbol{q}_{l}$; accordingly, it does not need to be re-computed when running (9).
Remark III.3 (Validity of the linear approximation).
The linear model is based on the bus admittance matrix $\boldsymbol{Y}$ and has constant matrices. This approximation is accurate for lightly loaded systems [41]. For heavily loaded system, [42, 10] showed that feedback-based methods are robust against model mismatch because of the closed-loop implementation; this feature is also pointed out in [34], and our analysis in the next section will characterize this robustness. The bus admittance matrix may be hard to obtain because it requires knowledge of the feeder characteristics, i.e., the line impedances, and the network configuration. However, one can assume that the system operator can obtain some estimates. Furthermore, in the case of network reconfiguration, the bus admittance matrix changes, leading to an incorrect linear model. However, this is not a frequent event and the system operator can update its linear model approximation when a reconfiguration occurs. $\Box$
In the next section, we analyze the convergence and stability properties of the proposed feedback-based SGF (9)
III-C Stability analysis and constraint satisfaction guarantees
@@
Lemma III.5 (Lipschitz continuity).
@@
Lemma III.5 follows from [43, Theorem 3.6], and by the compactness of the sets $\mathcal{E}_{M}$ and $\mathcal{E}_{J}$. This result ensures existence and uniqueness of solutions for the proposed feedback-based safe gradient flow [44, Ch. 3].
@@
Theorem III.6 (Practical local exponential stability).
@@
Lemma III.7 (Practical forward invariance).
Let the conditions in Theorem III.6 be satisfied, and let $\boldsymbol{u}(t)$, $t\geq t_{0}$, be the unique trajectory of (9), and $\boldsymbol{\nu}(t)$ be the corresponding voltage magnitudes. Define the set
@@
The proof is provided in Appendix -C. Lemma III.7 establishes forward invariance of a set $\mathcal{F}_{s}$, which is a subset of $\mathcal{F}_{e}$ and an inflation of $\mathcal{F}$ (more details about $\mathcal{F}_{s}$ are provided in the proof); clearly, $\mathcal{F}_{e}$ tends to the set $\mathcal{F}$ with the decreasing of the error in the computation of the Jacobian and the measurement errors, which implies that $\mathcal{F}_{s}$ tends to $\mathcal{F}$ too. If these errors are small, the voltage violation is practically negligible.
@@
Corollary III.8 (Error-free implementation).
Let all the conditions in Theorem III.6 be satisfied, and assume that there are no measurement errors, i.e., $E_{M}=0$, and no errors in the Jacobian, i.e., $E_{J}=0$ and. Let $\boldsymbol{u}(t)$, $t\geq t_{0}$, be the unique trajectory of (7). Then it holds that
@@
Lemma III.9 (Forward invariance in error-free implementation).
Let the conditions in Theorem III.6 be satisfied, and assume that there are no measurement errors (i.e., $E_{M}=0$). Let $\boldsymbol{u}(t)$, $t\geq t_{0}$, be the unique trajectory of (7), and $\boldsymbol{\nu}(t)$ be the corresponding trajectory of the voltages. Then, (7) renders the set $\mathcal{F}$ forward-invariant. In particular: (i) if $\nu_{i}(t_{0})\in[\underline{V},\bar{V}]$, then $\nu_{i}(t)\in[\underline{V},\bar{V}]$ for all $t\geq t_{0}$; (ii) if $\nu_{i}(t_{0})\notin[\underline{V},\bar{V}]$, then there exists $t^{\prime}\geq t_{0}$ such that $\nu_{i}(t)\in[\underline{V},\bar{V}]$ for all $t\geq t^{\prime}$. $\Box$
Corollary III.8 quantifies the error in the convergence to $\boldsymbol{u}^{*}$, and certify local exponential stability properties for the proposed method. Lemma III.9 establishes that, for the case with no measurement errors and with the exact computation of the Jacobian matrix, the proposed method ensures that voltages are satisfied anytime.
III-D Extensions
Virtual power plants. In this section, we introduce an extension of our formulation for virtual power plants (VPPs). In this case, the goal is to coordinate the operation of the DERs to regulate voltages and to provide ancillary services to the bulk power system. In particular, the coordination is to ensure that the active and reactive powers at the substation track a reference $\{P_{0,set},Q_{0,set}\}$. The reference setpoint can be sent by the transmission system operator in order to provide frequency regulation or ancillary services.
@@ III-D Extensions
Other constraints. The AC OPF formulation (4) can be extended to include constraints on the power flows or currents. The proposed feedback-based safe gradient flow can be naturally modified to enforce constraints on power flows or currents; to this end, one can use a linear model approximating the relationship between power injections at the DERs’ nodes and power flows and currents in the controller design [16]. During the operational phase, the controller would rely on measurements of power flows and currents (or pseudo-measurements) in addition to voltages; see, for example, the discussion in [16].
@@ IV-A Simulation setup
It is assumed that $p_{\text{max},i}$ is known at the DERs. For example, one can use the method proposed in [47] to estimate the maximum power point of PV arrays, and therefore, the maximum available power $p_{\text{max},i}$. Finally, we consider the following cost function for the SGF, PDM, and BO:
$C_{P}(p_{i},q_{i})=\sum_{i\in{\mathcal{G}}}c_{p}\left(\frac{s_{n,i}-p_{i}}{s_{n,i}}\right)^{2}+c_{q}\left(\frac{q_{i}}{s_{n,i}}\right)^{2}$ (28)
with $c_{p}=3$ and $c_{q}=1$. This cost function seeks to minimize active power curtailment and inverter power losses. The first part minimizes the active power curtailment, the second part the reactive power usage, which is also related to the inverter losses as less reactive power usage means less currents and thus, less Joules losses.
Volt/Var control
@@
Figure 4: (a) Overvoltages for NC. (b) Overvoltages for VVC. (c) Overvoltages for PDM. (d) Overvoltages for SGF. (e) Overvoltage duration times. (f) Linear map error: $||\hat{H}(\boldsymbol{u};\boldsymbol{p}_{l},\boldsymbol{q}_{l})-H(\boldsymbol{u};\boldsymbol{p}_{l},\boldsymbol{q}_{l})||$ and Jacobian error: $||J_{\hat{H}}-J_{H}(\boldsymbol{u};\boldsymbol{p}_{l},\boldsymbol{q}_{l})||$, where $\boldsymbol{u}$ is picked from the SGF algorithm.
@@ IV-B Results
In Fig. 4, one can see the maximum voltage observed at every time step, as well as the number of impacted nodes where we observed a voltage greater than $\bar{V}$. It can be seen that only the SGF method does not lead to voltage violations. This is precisely because our approach is based on the theory of control barrier function. PDM leads to voltage violations not only because of the transient of the dual variables, but also because it is designed based on a regularization of the Lagrangian function as explained in [10]. VVC performs well, although as shown hereafter, it leads to larger system losses and a greater cumulative cost than SGF or PDM. The overall voltage profile is also shifted downward due to its proportional feedback control. One can observe the spikes in the NC method due to multiple disconnections of DERs because of a prolonged overvoltage duration. Finally, one can see that with PDM, the voltages oscillate around the threshold voltage of 1.05 pu.
@@
Finally, Fig. 4(f) shows the error between the linear approximation of the power flow equations and the non-linear power flow equations, validating our choice for the linear map. It also shows the error between the approximate Jacobian (which is constant), and the true Jacobian computed numerically.
@@ V Conclusions
By [25, Lemma 5.11 and Theorem 5.6(iii)], $\bar{F}_{\beta}(\boldsymbol{u},\mymathbb{0},\mymathbb{0})$ is differentiable at $\boldsymbol{u}^{*}$ and its Jacobian $E=\frac{\partial\bar{F}_{\beta}(\boldsymbol{u},\mymathbb{0},\mymathbb{0})}{\partial\boldsymbol{u}}\mid_{\boldsymbol{u}=\boldsymbol{u}^{*}}$ is negative definite. Recall that $e_{1}=-\lambda_{\max}(E)$ and $e_{2}=-\lambda_{\min}(E)$. Let $P:=\int_{0}^{\infty}(\exp(E\zeta)^{\top}\exp(E\zeta)d\zeta$, and then by [44, Theorem 4.12], it holds that $PE+E^{\top}P=-\mathbf{I}_{n}$, and $\frac{1}{2e_{2}}\|u-u^{*}\|_{2}^{2}\leq(u-u^{*})^{\top}P(u-u^{*})\leq\frac{1}{2e_{1}}\|u-u^{*}\|_{2}^{2}$. Let $V_{1}(\boldsymbol{u}):=$$(\boldsymbol{u}-\boldsymbol{u}^{*})^{\top}P(\boldsymbol{u}-\boldsymbol{u}^{*})$; then we bound $2(\boldsymbol{u}-\boldsymbol{u}^{*})^{\top}P\bar{F}_{\beta}(\boldsymbol{u},\mymathbb{0},\mymathbb{0})$ and then leverage this bound to estimate $\dot{V}_{1}$:
@@
In addition, we note that that for any $a\geq 0,~{}b>0$, $y(t)=y(t_{0})\exp{(-b(t-t_{0}))}+\frac{a}{b}(1-\exp{(-b(t-t_{0})})$ is the solution of $\dot{y}=-by+a,y(t_{0})=y(t_{0})$. Hence by the Comparison Lemma [44, Lemma 3.4], it follows that